OpenStudy (anonymous):

Solve. 6 – 3x = 4(3 – x) – 6

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

6 – 3x = 4(3 – x) – 6 6 – 3x =12 – 4x – 6 6 – 3x =6 – 4x 4x-3x= 6-6 x=0 @zobobozo

OpenStudy (anonymous):

0